# Topological properties of curved spacetime Su–Schrieffer–Heeger model

### v2

@@ VIII Summary and Discussion
In this paper, we investigate the topological properties of the CST-SSH model. First, we analyze the energy spectrum and find that when the intercell hopping is larger than the intracell hopping, a pair of zero-energy states emerges. These states are localized at the two edges of the lattice. However, unlike in the conventional SSH model, these edge states are not symmetric. We also find that, similar to the usual SSH model, the spectrum exhibits a gap whenever the intra- and inter-cell hopping amplitudes are unequal. However, this gap scales to zero with system size as $N^{-\sigma}$. Since the typical gap of a finite-size system scales as $N^{-1}$, it implies that for $\sigma<1$ the spectrum remains gapped in the thermodynamic limit [49]. Moreover, we use various real-space topological markers, both static and dynamic, which show a clear signature of a topological phase transition between a topologically trivial phase and a non-trivial phase with winding number $1$. On the other hand, we find that the physics of the event horizon manifests only at the transition point, where a critical slowdown occurs for zero-energy wave packets near the boundary. The wave-packet dynamics results are also supported by an analytical calculation of semiclassical trajectories. We have also generalized our results to the extended CST-SSH model, which can support topological phases with higher winding numbers. In this model as well, we demonstrate that the critical slowdown of the wavepacket occurs only at the transition point between two topologically distinct phases. With the advancement of cold-atom experiments, there is potential for our predictions to be verified in the near future [58, 59, 60], which could enrich our understanding of both black hole physics and topology.
@@ Acknowledgements.
Appendix A Zero edge states of the Extended CST-SSH Model
Figure 12: Closing of the energy band gap and variation of zero energy eigen states variation with $\sigma$: for the extended CST-SSH model having system size $2N=1000$ (500 unit cells) with $(t_{1},t_{2},t_{3},t_{4})=(2,1,-1,-4)$, which have a winding number 2. Top row (a),(b): Eigenvalue spectra for $\sigma$=0 and 1, respectively. Bottom row (c),(d): Corresponding zero-energy eigenstates profiles.
In this section, we discuss both the energy band closing and the zero-energy edge states in the extended version of the CST-SSH model. For the Hamiltonain  5, we observe that various winding numbers—specifically 1-1, 0, 11, and 22—can be realized by appropriately tuning the parameters $(t_{1},t_{2},t_{3},t_{4})$. Here, the winding number corresponds to the number of zero-energy edge modes per boundary under open boundary conditions.
For example, in Fig. 12, we choose $(t_{1},t_{2},t_{3},t_{4})=(2,1,-1,-4)$, which yields a winding number of two. This is evident from the presence of two zero-energy edge states per boundary, observed for both $\sigma=0$ and $\sigma=1$. The model can be further extended or modified to realize higher winding numbers.
Similar to the CST-SSH model discussed in the main text, the energy gap in the extended model gradually closes as a function of $\sigma$, and the zero-energy edge states become less localized. However, the application of $\sigma$ does not alter the winding number, which was demonstrated in the Fig.10.

### v3

@@ VIII Summary and Discussion
In this paper, we investigate the topological properties of both the extended and standard version of CST-SSH model. We find that, for both model, the physics of the event horizon manifests only at the topological phase transition points, where a critical slowdown occurs for zero-energy wave packets near the boundary. The wave-packet dynamics results are also supported by an analytical calculation of semiclassical trajectories. We analyze the energy spectrum and find that when the intercell hopping is larger than the intracell hopping, a pair of zero-energy states emerges. These states are localized at the two edges of the lattice. However, unlike in the conventional SSH model, these edge states are not symmetric. We also find that, similar to the usual SSH model, the spectrum exhibits a gap whenever the intra- and inter-cell hopping amplitudes are unequal. However, this gap scales to zero with system size as $N^{-\sigma}$. Since the typical gap of a finite-size system scales as $N^{-1}$, it implies that for $\sigma<1$ the spectrum remains gapped in the thermodynamic limit [36], and gapless for $\sigma\geq 1$. Moreover, we use various real-space topological markers, both static and dynamic, which show a clear signature of a topological phase transition for the CST-SSH models between a topologically trivial phase and a non-trivial phase for the exact parameters, as the usual SSH models. Even though the spectrum remains gapless away from the transition points for $\sigma\geq 1$, these regions do not exhibit horizon physics. This shows that the mere gaplessness of the spectrum is not sufficient for the emergence of horizon physics. Instead, the gapless points corresponding to topological transitions are special, as they support the critical slowdown associated with horizon physics. This observation naturally raises a fundamental question: how can one define a topological marker that characterizes transitions between phases that remain gapless in the CST versions of SSH-type models? Despite the introduction of power law position dependent hopping, we show that the symmetry class of the CST-SSH models remain in BDI category. Moreover, it turns out that although the spectrum is gapless, the states in the neighborhood of the zero-energy modes are spatially localized (see Appendix A). This localization is likely the reason why the system retains its topological character. Effectively, a mobility gap emerges in the system. In other words, the zero-energy states are surrounded by localized states that do not contribute to transport, making the system effectively insulating. Such phenomena are typically observed in disordered topological systems [64, 62]. Even very recently, such phenomena have been reported in disordered SSH models, where, in contrast to the conventional SSH model, the topological transition points separate two gapless phases [37]. We also study the quench dynamics from topological to trivial state and explain the oscillatory behavior of the survival probability with our semiclassical calculations. Further, analyze the slowdown quantification for the extended CST-SSH, and found that, in the vicinity of the higher order gap closing points, where two gap closing conditions for the conventional extended SSH models are satisfied for the same $k_{c}$ value, the slowdown rate is higher. Also, for the other higher order gap closing points, where two gap closing conditions are satisfied for two different $k_{c}$ values, we observe that there are two distinct values of $p_{0}$ $(=\pi,\pi/2)$, having two critical slowdown with different $\tau_{s}$, for the same set of hopping parameters. With the advancement of cold-atom experiments, there is potential for our predictions to be verified in the near future [63, 58, 3], which could enrich our understanding of both black hole physics and inhomogeneous topological systems.
@@ Acknowledgements.
Appendix A Zero energy and its neighboring states of the extended CST-SSH Model
[…]
Figure 13: IPR of the eigen states of the near zero energy neighborhodd, for $\sigma=0,1$ for different system sizes with $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$,
In this section, we discuss both the energy band closing and the zero-energy edge states in the extended CST-SSH model. For the Hamiltonian  4, we observe various winding numbers: specifically 1-1, 0, 11, and $2$, which can be realized by appropriately tuning the parameters $(t_{1},t_{2},t_{3},t_{4})$. Here, the winding number corresponds to the number of zero-energy edge modes per boundary under open boundary conditions.
For example, in Fig. 12, we choose $(t_{1},t_{2},t_{3},t_{4})=(2,1,-1,-4)$, which yields a winding number of two. This is evident from the presence of two zero-energy edge states per boundary, observed for both $\sigma=0$ and $\sigma=1$. Figs. 12e, 12f shows the probability densities corresponding to two states near to the zero energy eigen states. For $\sigma=0$, they are delocalized; but at $\sigma=1$, these states become localized near the edge. The model can be further extended or modified to realize higher winding numbers.
As discussed in the main text, the energy gap in the extended model gradually closes as a function of $\sigma$, and one of the zero-energy edge states become more localized. However, the application of $\sigma$ does not alter the winding number, which was demonstrated in the Fig.6. We also study the wavefunctions of the states in the neighborhood of the zero-energy localized states (Fig. 12f) and find that, for the CST-SSH model, the probability densities corresponding to these states are also spatially localized. To probe more into these localized states, we also calculated the Inverse Participation ratio (IPR) for the states in the neighborhood of the zero energy states. To quantify the degree of localization of an eigenstate $|\psi\rangle=\sum_{i}\psi_{i}\,|i\rangle$, we compute the inverse participation ratio (IPR), defined as [21]
$\text{IPR}=\sum_{i}|\psi_{i}|^{4},$ (26)
for which larger values indicate stronger spatial localization. The Fig. 13 shows the IPR for various system sizes and for $\sigma=0,1$. For $\sigma=0$, IPR decreases as we increase the system size (except the two zero energy states at the centre), implying the eigenstates are delocalized (except the zero energy states). For $\sigma=1$, the states around zero energy, with increasing the system size, have the same IPR values, thus confirming that these zero energy and even their neighborhood states are localized. As a result, these states do not contribute to transport and effectively act as a mobility gap, even though the spectrum itself remains gapless. This behavior is absent in the $\sigma=0$ case.